integration of  ((|x|)/([x]) from (1,9)..

Grade:

1 Answers

Chetan Mandayam Nayakar
13 years ago

throughout the range of x, mod(x)=x, I=∫(x/[x])dx from 1 to 9, =∫(x/[x])dx from 1 to 2...from 8 to 9

∑(((i+1)2-i2)/2)/i    from i=1 to i=8

=3/2 + 5/4 +7/6 +9/8 = 121/24

'Win exciting gifts by answering the questions on Discussion Forum.'